# General effective reduction theory of integral polynomials of given non-zero discriminant and its applications

### v1

@@ 4.1. The theorems of Birch and Merriman (1972), Győry (1973) and Evertse and Győry (1991)
For monic polynomials, the corresponding result with $\mathbb{Z}$-equivalence was proved independently by Győry (1973) but in an effective form. This proved to be of crucial importance in many applications; see e.g. Sections 5 to 8 below and Evertse and Győry (2017).
@@
where $c_{1}$ and $c_{2}$ are effectively computable positive numbers depending on $D$, resp. on $n$ and $|D|$.
@@
There are only finitely many $\mathbb{Z}$-equivalence classes of monic polynomials in $\mathbb{Z}[X]$ of given non-zero discriminant, and a full set of representatives of these classes can be effectively determined.
@@
Theorem 4.1, resp. Theorem 4.2 and its Corollary 4.3 generalized the corresponding results presented in Section 2 to polynomials of any degree $n\geq 3$; Theorem 4.1 gives an ineffective generalization and Theorem 4.2 an effective generalization.
@@
Let $f\in\mathbb{Z}[X]$ be a polynomial of degree $n\geq 2$ and discriminant $D\neq 0$. There is a $g\in\mathbb{Z}[X]$, $GL_{2}(\mathbb{Z})$-equivalent to $f$, such that
@@
As was mentioned above, Theorems 4.2 and 4.4 provided an effective version of the general reduction theory of integral polynomials of given non-zero discriminant.
@@
Theorems 4.2 and 4.4, their explicit versions below and their various generalizations have a great number of consequences and applications; see our book Evertse and Győry (2017) and Sections 5 to 8 below.
@@ 4.2. Explicit versions of theorems of Győry (1973) and Evertse and Győry (1991a)
As was mentioned above, the first explicit version of Theorem 4.2 was given in Győry (1974). This was improved in 2017 by the authors.
@@
(4.1) $H(g)\leq\exp\{n^{20}8^{n^{2}+19}(|D|(\log|D|)^{n})^{n-1}\}.$
@@
Corollary 4.6 (Evertse and Győry, 1991).
There are only finitely many $GL_{2}(\mathbb{Z})$-equivalence classes of polynomials in $\mathbb{Z}[X]$ of given non-zero discriminant, and a full set of representatives of these classes can be effectively determined.
$c_{4}(n,G)|D|^{c_{5}(n,G)},$
where $c_{4}(n,G),c_{5}(n,G)$ are effectively computable numbers which depend only on $n$ and the discriminant of $G$; see Győry (1984, 1998) resp. Evertse and Győry (1991a). The following conjecture seems plausible.
Conjecture 4.7.
Let $f\in\mathbb{Z}[X]$ be a polynomial resp. a monic polynomial of degree $n\geq 4$ resp. $n\geq 3$ with discriminand $D\neq 0$. Then $f$ is $GL_{2}(\mathbb{Z})$-equivalent resp. $\mathbb{Z}$-equivalent to a polynomial resp. monic polynomial such that
$H(g)\leq c_{6}(n)|D|^{c_{7}(n)}$
where $c_{6}(n),c_{7}(n)$ depend only on $n$.
The first part of the conjecture is formulated in Section 15 of Evertse and Győry (2017).
Evertse proved the following what one may call semi-effective result.
Theorem 4.8 (Evertse, 1993).
$H(g)\leq c_{8}(n,G)|D|^{21/n}.$
Here $c_{8}(n,G)$ is a number depending only on $n$ and $G$, which is not effectively computable by the method of proof. Evertse’ s proof is based on a version of Roth’s Diophantine approximation theorem over number fields, which is ineffective.
4.3. The general approach of Győry (1973) for proving Theorem 4.2
@@
For the complete proof we refer to Győry (1973) or the Subsections 6.6 and 6.7 of Evertse and Győry (2017). Here, we give a brief outline. The primitive minimal polynomial $f_{\alpha}$ of an algebraic number $\alpha$ is the minimal polynomial of $\alpha$ whose coefficients are integers with greatest common divisor $1$ and whose leading coefficient is positive. The height $H(\alpha)$ of an algebraic integer $\alpha$ is the height of $f_{\alpha}$.
The proof of the boundedness of the degree $n$ of $f$ in terms of $D$ uses Minkowski’s lower bound for the absolute value of the discriminant of a number field in terms of its degree and combines this with an elementary argument. Henceforth, we restrict ourselves to the case that $f\in\mathbb{Z}[X]$ is a monic irreducible polynomial of discriminant $D\neq 0$ and of fixed degree $n$ with $2\leq n\leq c_{1}(|D|)$.
Thus, let $f\in\mathbb{Z}[X]$ be such a polynomial. Then the main steps of the proof of Theorem 4.2 are as follows.
  1. (1)

### v2

@@ 4.1. The theorems of Birch and Merriman (1972), Győry (1973) and Evertse and Győry (1991a)
For monic polynomials, the corresponding result with $\mathbb{Z}$-equivalence was proved independently by Győry (1973) but in an effective form. This turned out to be of crucial importance in many applications; see e.g. Sections 5 to 9 below and Evertse and Győry (2017).
@@
where $c_{1}$ and $c_{2}$ are effectively computable positive numbers depending on $|D|$, resp. on $n$ and $|D|$.
This theorem was first proved and published in Győry’s PhD dissertation Győry (1972a) and was utilized in Győry (1972b) as well.
@@
There are only finitely many $\mathbb{Z}$-equivalence classes of monic polynomials in $\mathbb{Z}[X]$ of given non-zero discriminant, and a full set of representatives of these classes can be at least in principle determined.
@@
Theorem 4.1, resp. Theorem 4.2 and its Corollary 4.3 are generalizations of the corresponding results presented in Section 2 for polynomials of degree $n\leq 3$; Theorem 4.1 gives an ineffective generalization of Theorem 2.4 for degree $n\geq 4$ and Theorem 4.2 is an effective generalization of Theorem 2.3 in the monic case for degree $n\geq 3$, and of Theorem 2.5 for any monic polynomial of degree $n\geq 3$.
@@
Let $f\in\mathbb{Z}[X]$ be a polynomial of degree $n\geq 2$ and discriminant $D\neq 0$. There is $g\in\mathbb{Z}[X]$, $GL_{2}(\mathbb{Z})$-equivalent to $f$, such that
@@
This theorem was stated and proved in Evertse and Győry (1991a) in an equivalent form, in terms of integral binary forms.
As was mentioned above, Theorems 4.2 and 4.4 led to a general effective reduction theory of integral polynomials of given non-zero discriminant.
@@
We note that Theorems 4.1 and 4.4 were established directly in a more general form, in the number field and $\mathfrak{p}$-adic case. For such and other generalizations of Theorem 4.2, (ii), see Győry (1978b, 1984) and Section 9 below.
Theorems 4.2 and 4.4, their explicit versions below and their various generalizations have a great number of consequences and applications; see our book Evertse and Győry (2017) and Sections 5 to 9 below.
@@ 4.2. Explicit versions of theorems of Győry (1973) and Evertse and Győry (1991a)
As was mentioned above, the first explicit version of Theorem 4.2 was given in Győry (1974). The height estimate was improved in 2017 by the authors.
We use the notation $\log^{\ast}x:=\max(1,\log x)$ for $x>0$.
@@
(4.1) $H(g)\leq\exp\{n^{20}8^{n^{2}+19}(|D|(\log^{\ast}|D|)^{n})^{n-1}\}.$
This is in fact Theorem 6.6.2 from Evertse and Győry (2017) with a slightly larger, simplified constant in terms of $n$.
@@
Corollary 4.6 (Evertse and Győry, 1991a).
There are only finitely many $GL_{2}(\mathbb{Z})$-equivalence classes of polynomials in $\mathbb{Z}[X]$ of given non-zero discriminant, and a full set of representatives of these classes can be at least in principle effectively determined.
4.3. Consequences of Theorems 4.4*, 4.2* and Theorem 3.3 for Hermite equivalence classes
@@
As was pointed out in BEGyRS (2023),an important consequence of the above Theorem 3.3 is that the effective finiteness theorems 4.4, 4.4* and 4.2, 4.2* for $GL_{2}(\mathbb{Z})$-equivalence classes resp. $\mathbb{Z}$-equivalence classes apply just as well to Hermite equivalence classes.
We present here the following, more precise, explicit variant of Hermite’s result in Theorem 3.2.
Corollary 4.7 (of Theorems 4.4* and 4.2*; cf. BEGyRS, 2023).
  1. (i)
    Every Hermite equivalence class of polynomials in $\mathbb{Z}[X]$ of degree $n\geq 2$ and of discriminant $D\neq 0$ has a representative with coefficients not exceeding
    $\exp\{(4^{2}n^{3})^{25n^{2}}|D|^{5n-3}\}$
    in absolute value.
  2. (ii)
    Every Hermite equivalence class of monic polynomials in $\mathbb{Z}[X]$ with degree $n\geq 2$ and discriminant $D\neq 0$ has a representative with coefficients not exceeding
    $\exp\{n^{20}8^{n^{2}+19}(|D|(\log^{\ast}|D|)^{n})^{n-1}\}$
    in absolute value.
It is an immediate consequence of Theorem 4.5 that in (i) above $n\leq 3+2\log|D|/\log 3$. Further, in (ii), the slightly better inequality $n\leq 2+2\log|D|/\log 3$ holds.
The above result implies an effective version of Theorem 3.2, i.e., for given $n$ and a non-zero integer $D$, one can effectively determine a full system of representatives for the Hermite equivalence classes of polynomials $f\in\mathbb{Z}[X]$ of degree $n$ and discriminant $D$. Indeed, one can make a finite list of all polynomials $f\in\mathbb{Z}[X]$ of height below one of the bounds in Corollary 4.7. For each polynomial in the list one can check whether it has discriminant $D$. Further, for each pair of polynomials in the list one can check whether they are Hermite equivalent, by computing the corresponding decomposable forms $[f],[g]$, and checking whether they are $GL_{n}(\mathbb{Z})$-equivalent, using, e.g., Lemma 18 of Evertse and Győry (1992a).
The similarity of Theorems 4.4* , 4.2* and Corollary 4.7 is only apparent. As was seen in Section 3, the $GL_{2}(\mathbb{Z})$-equivalence and $\mathbb{Z}$-equivalence are in fact much stronger than the Hermite equivalence.
Remark.
Every improvement of the bounds in (4.1) or (4.2) would yield the same improvement in the bounds of Corollary 4.7.
4.4. Unit equations and $S$-unit equations
a
The unit equations and more general $S$-unit equations play a fundamental role in Diophantine number theory, and in particular in the effective reduction theory of integral polynomials of given discriminant.
First we recall unit equations and $S$-unit equations, and then briefly outline how to apply Baker’s theory of logarithmic forms to obtain effective bonds for the solutions of these equations. Then we recall the best known height bounds for the solutions of unit equations and $S$-unit equations over number fields.
For a detailed treatment of unit equations, $S$-unit equations and their further generalizations and applications we refer to our books Evertse and Győry (2015, 2017, 2022).
Let $K$ be an algebraic number field, $\mathcal{O}_{K}$ its ring of integers, $\mathcal{O}_{K}^{\ast}$ the unit group of $\mathcal{O}_{K}$, and $M_{K}$ its set of places, consisting of the finite set of infinite places $S_{\infty}$ of $K$ (corresponding to the real embeddings and the pairs of conjugate complex embeddings of $K$ in $\mathbb{C}$) and the finite places, which we may identify with the prime ideals of $\mathcal{O}_{K}$. To the places in $M_{K}$ we can associate a set of absolute values $\{|\cdot|_{v}:\,v\in M_{K}\}$, normalized such that if $v$ lies above the place $p\in M_{\mathbb{Q}}:=\{\infty\}\cup\{\text{primes}\}$, then for $a\in\mathbb{Q}$ one has $|a|_{v}=|a|_{p}^{[K_{v}:\mathbb{Q}_{p}]}$. These absolute values satisfy the product formula $\prod_{v\in M_{K}}|\alpha|_{v}=1$ for $\alpha\in K^{*}$.
Let $a,b$ be given non-zero elements of $K$. Equations of the form
(4.3) $ax+by=1\text{ in unknowns }x,y\in\mathcal{O}_{K}^{\ast}$
are called unit equations (in two unknowns). More generally, let $S$ be a finite subset of $M_{K}$ with $S\supseteq S_{\infty}$. Denote by $\mathcal{O}_{S}$ the ring of $S$-integers, i.e., $\{x\in K:\,|x|_{v}\leq 1\ \text{for }v\not\in S\}$ and by $\mathcal{O}_{S}^{\ast}$ denote the unit group of $\mathcal{O}_{S}$, i.e., group of $S$-units. Thus, $\mathcal{O}_{S}^{\ast}=\{x\in K:\,|x|_{v}=1\ \text{for }v\not\in S\}$. For $S=S_{\infty}$ we have $\mathcal{O}_{S}^{\ast}=\mathcal{O}_{K}^{\ast}$. Equations of the form
(4.4) $ax+by=1\text{ in unknowns }x,y\in\mathcal{O}_{S}^{\ast}$
are called $S$-unit equations (in two unknowns). In many cases it is more convenient to consider the unit equations and $S$-unit equations in homogeneous form
(4.4a) $ax+by+cz=0\text{ in unknowns }x,y,z\in\mathcal{O}_{K}^{\ast},\ \text{resp. }\mathcal{O}_{S}^{\ast},$
where $a,b,c$ denote fixed elements of $K\setminus\{0\}$.
For a long time these equations were utilized merely in special cases and in an implicit way. It was implicitly proved by Siegel (1921) for $S=S_{\infty}$ and by Parry (1950) for any $S$ that equation (4.4) has only finitely many solutions. This implies the finiteness of the number of solutions of equation (4.4a) up to a common proportional factor. Lang (1960) gave a direct proof for a more general version of these finiteness theorems. Their proofs were ineffective.
Generalizing Gelfond’s (1935) famous result obtained in the case $m=2$, in the 1960’s Baker made a major breakthrough in number theory by giving non-trivial explicit lower bounds for the absolute value of linear forms in logarithms of the form
$b_{1}\log\alpha_{1}+\cdots+b_{m}\log\alpha_{m}\neq 0,\,\,m\geq 2$
where $b_{1},\ldots,b_{m}$ are rational integers, resp. algebraic numbers, $\alpha_{1},\ldots,\alpha_{m}$ are algebraic numbers different from 0 and $1$, and $\log\alpha_{1},\ldots,\log\alpha_{m}$ denote fixed determination of the logarithms. In case of rational integers $b_{1},\ldots,b_{m}$, this is equivalent to bounding $|\prod\alpha_{i}^{b_{i}}-1|$ non-trivially from below. Baker’s general effective estimates led to significant applications, and opened a new effective epoch in the theory of Diophantine equations. Baker’s quantitative results were later improved, generalized, extended to the $\mathfrak{p}$-adic case and so on by himself and many other authors; for comprehensive overviews we refer to Baker (1990), Wüstholz, ed. (2002), and Baker and Wüstholz (2007), and for a shorter overview see Evertse and Győry (2015), Section 3.2. The last five decades saw the development of an effective theory of Diophantine equations.
General effective upper bounds for the solutions of (4.3) and (4.4a) in the case $S=S_{\infty}$ were deduced by Győry (1972a,b, 1973) using an effective result of Baker and Coates (1970), p. 601, on relative Thue equations over number fields. The first explicit upper bounds for the solutions of (4.3) and (4.4a) in case $S=S_{\infty}$ were deduced by Győry (1974) from an explicit inequality of Baker (1968a, Part IV) for linear forms in logarithms of algebraic numbers. For general $S$, Győry (1979) derived the first explicit bound for the solutions of (4.4), using also the $\mathfrak{p}$-adic version of Baker’s theory. Independently, a slightly weaker effective bound was given by Kotov and Trelina (1979).
Let $K$ be an algebraic number field. Given $\alpha_{1},\ldots,\alpha_{n}\in K$, not all 0, we define the height of $(\alpha_{1},\ldots,\alpha_{n})$ relative to $K$ by
$H_{K}(\alpha_{1},\ldots,\alpha_{n}):=\prod_{v\in K}\max(|\alpha_{1}|_{v},\ldots,|\alpha_{n}|_{v}).$
Recall that the naive height $H(\alpha)$ of an algebraic number $\alpha$ is given by the maximum of the absolute values of its minimal polynomial, with coefficients having gcd $1$. Then we have
$H(\alpha)\leq 2^{\deg\alpha}H_{\mathbb{Q}(\alpha)}(1,\alpha).$
Following Section 1.3 from the paper "Solving Diophantine equations by Baker’s theory" by Győry (2002), we briefly sketch a proof of the following theorem, by means of Baker’s theory.
Theorem 4.8.
Let $K$ be a number field, $S$ a finite set of places of $K$ containing $S_{\infty}$, and $a,b$ non-zero elements of $K$. Let $x,y\in\mathcal{O}_{S}^{\ast}$ satisfy (4.4). Then
$\max(H(x),H(y))\leq c_{4}(K,S,a,b),$
where $c_{4}$ is an effectively computable number, depending only on $K,S,a,b$.
Sketch.
Let $s$ denote the cardinality of $S$. There is a system of fundamental $S$-units $\{\varrho_{1},\ldots,\varrho_{s-1}\}$ in $\mathcal{O}_{S}^{*}$ with heights bounded in terms of $K$ and $S$. Let $x,y$ be a solution of (4.4) in $S$-units. Then one can write
$x=\xi_{1}\varrho_{1}^{a_{11}}\cdots\varrho_{s-1}^{a_{1,s-1}},\,\,y=\xi_{2}\varrho_{1}^{a_{21}}\cdots\varrho_{s-1}^{a_{2,s-1}},$
where $\xi_{1},\xi_{2}$ are roots of unity in $K$ and $a_{ij}$ are unknown rational integer exponents. Assume without loss of generality that $A:=\max_{j}|a_{1j}|\geq\max_{j}|a_{2j}|$. By elementary means one can show that
$A\leq c_{5}\log\max_{v\in S}|x|_{v},$
and combining this with $\prod_{v\in S}|x|_{v}=1$, one concludes that there is a $v\in S$ such that
$|x|_{v}\leq c_{6}\exp\{-c_{7}A\},$
where $c_{5},c_{6},c_{7}$ can be given explicitly and depend only on $K$ and $S$. This implies
(4.5) $0<|\varrho_{1}^{a_{21}}\cdots\varrho_{s-1}^{a_{2,s-1}}-\alpha|_{v}\leq c_{8}\exp\{-c_{9}A\}$
with an appropriate $\alpha\in K$ of bounded height. The constants $c_{8},c_{9}$ and $c_{10}$ below depend at most on $K,S$ and $a,b$ and can be given explicitly.
One can now apply the complex or $\mathfrak{p}$-adic version of Baker’s theory according as $v\in S_{\infty}$ or $v\in S\setminus S_{\infty}$ and this yields
$\exp\{-c_{10}\log A\}\leq|\varrho_{1}^{a_{21}}\cdots\varrho_{s-1}^{a_{2,s-1}}-\alpha|_{v}.$
Comparing this with (4.5) we get
(4.6) $A\leq A_{0}$
where $A_{0}$ can be given explicitly. Finally, we obtain an upper bound for $H(x)$ and $H(y)$ which can also be given explicitly. ∎
Later, several improvements, effective generalizations, applications and algorithmic results have been obtained for unit and $S$-unit equations by means of Baker’s theory; see among others Győry (1980b, 2002, 2019, 2022), Shorey and Tijdeman (1986), Sprindžuk (1993), Bugeaud and Győry (1996), Smart (1998), Gaál and Győry (1999), Hindry and Silverman (2000), Wüstholz, ed. (2002), Bilu (2002), Bilu, Gaál and Győry (2004), Győry and Yu (2006), Baker and Wüstholz (2007), Zannier (2009), Hajdu (2009), Bérczes, Evertse and Győry (2009), Evertse and Győry (2013, 2015 2017, 2022), Bérczes (2015a, 2015b), Bertók and Hajdu (2015, 2018), Bugeaud (2018), Gaál (2019), Le Fourn (2020), Alvarado et al. (2021), Győry and Le Fourn (2024), and the references given there.
The best known height bound for the solutions of (4.3) is due to Győry and Yu (2006). We formulate it in simplified form. As above, let $K$ be a number field of degree $d$ and $r$ the rank of $\mathcal{O}_{K}^{*}$. Denote by $h_{K},\,R_{K}$ the class number and regulator of $K$, respectively, and write again $\log^{\ast}x:=\max(1,\log x)$.
Then Győry and Yu (2006) proved the following.
Theorem 4.9.
Let $a,b$ be non-zero elements of $K$. Then for all $x,y\in\mathcal{O}_{K}^{*}$ satisfying (4.3) we have
$H_{K}(1,x,y)\leq(3H_{K}(1,a,b))^{A},$
where
$A=d^{5}(2r+2)^{4r+40}R_{K}\log^{\ast}R_{K}.$
Remark. The following inequality implies that $A$ can be bounded abouve in terms of $d$ and $D_{K}$ only:
$h_{K}R_{K}\leq|D_{K}|^{1/2}(\log^{*}|D_{K}|)^{d-1}.$
The first inequality of this type was proved by Landau (1918). For the above version, see, e.g., Evertse and Győry (2015, formula (1.5.2)).
In this section we shall use only Theorem 4.9. Below we formulate a generalization to $S$-unit equations, Theorem 4.10, which is not used in this section, but will be needed in Section 9.
Consider the general case where $S$ is an arbitrary finite set of places containing $S_{\infty}$. In terms of $S$, the best known bounds can be found in Győry (2019), Le Fourn (2020) and Győry and Le Fourn (2024). We mention here the bound from Győry (2019) in simplified form. We introduce the necessary notation. Let as above $K$ be a number field of degree $d$ and $r$ the rank of $\mathcal{O}_{K}^{*}$. Denote by $h_{K}$, $R_{K}$ the class number and regulator of $K$, respectively. Further, let $S=S_{\infty}\cup\{\mathfrak{p}_{1},\ldots,\mathfrak{p}_{t}\}$, where $\mathfrak{p}_{1},\ldots,\mathfrak{p}_{t}$ with $t\geq 0$ are the prime ideals in $S$. Let s:=#Ss:=\#S, and denote by $R_{S}$ the $S$-regulator. It is known that
$R_{S}=R_{K}\ \text{if }t=0,\ \ R_{S}=i_{S}R_{K}\prod_{i=1}^{t}\log N_{K}\mathfrak{p}_{i}\ \text{otherwise,}$
where $i_{S}$ is a divisor of $h_{K}$ and $N_{K}\mathfrak{a}$ denotes the norm of a non-zero ideal $\mathfrak{a}$ of $\mathcal{O}_{K}$, i.e., $\mathcal{O}_{K}/\mathfrak{a}$. Let $P_{S}:=1$ if $t=0$ and $P_{S}:=\max_{1\leq i\leq t}N_{K}\mathfrak{p}_{i}$ if $t\geq 1$. Further, put $P_{S}^{\prime}:=1$ if $t\leq 2$ and $P_{S}^{\prime}$ the third largest among the quantities $N_{K}\mathfrak{p}_{i}$, $i=1,\ldots,t$ if $t\geq 3$. Finally, put $\mathcal{T}_{K}:=\max(h_{K},160r!\cdot(r+1)^{2}R_{K})$.
Theorem 4.10.
Let $a,b$ be non-zero elements of $K$. Then for all $x,y\in\mathcal{O}_{S}^{*}$ with (4.4) we have
$H_{K}(1,x,y)\leq(3H_{K}(1,a,b))^{A_{S}},$
where
$A_{S}:=2s^{5}(16ed)^{4s+3}\mathcal{T}_{K}^{t+4}\cdot\frac{P_{S}^{\prime}}{\log^{*}P_{S}^{\prime}}\Big(1+\frac{\log^{*}\log P_{S}}{\log^{*}P_{S}^{\prime}}\Big)R_{S}.$
Observe that for $S=S_{\infty}$, $A$ is much smaller than $A_{S}$. Further, $A_{S}$ can be bounded above in terms of $d$, $|D_{K}|$, $t$, and $P_{S}$.
We compare Theorem 4.10 with the abc-conjecture over number fields. We first recall the abc-conjecture over $\mathbb{Q}$, as proposed by Masser (1985), refining an earlier conjecture of Oesterlé. Define the radical of a non-zero integer $a$ by $\mathcal{R}(a):=\prod_{p|a}p$.
Conjecture 4.11 (Masser-Oesterlé abc-conjecture, 1985).
There is a constant $C(\epsilon)>0$ depending on $\epsilon$ such that for all $\epsilon>0$ and all non-zero integers $a,b,c$ with $a+b=c$ and ${\rm gcd}(a,b,c)=1$ we have $\max(|a|,|b|,|c|)\leq C(\epsilon)\mathcal{R}(abc)^{1+\epsilon}$.
There are various proposals to extend this to number fields. We recall a version of Masser (2002). Let $K$ be a number field and $D_{K}$ its discriminant. Take a non-zero ideal $\mathfrak{a}$ of $\mathcal{O}_{K}$. Masser defined the modified radical of $\mathfrak{a}$ by $\mathcal{R}_{K}(\mathfrak{a}):=\prod_{\mathfrak{p}|\mathfrak{a}}N_{K}\mathfrak{p}^{e_{\mathfrak{p}}}$, where the product is taken over all prime ideals dividing $\mathfrak{a}$ and $e_{\mathfrak{p}}$ is the ramification index of $\mathfrak{p}$. Masser considered this modified radical since it has a good behaviour under field extensions, e.g., if $L$ is an extension of $K$ of degree $m$, then $\mathcal{R}_{L}(\mathfrak{a}\mathcal{O}_{L})=\mathcal{R}_{K}(\mathfrak{a})^{m}$.
Recall that the different of $K$ can be expressed as $\mathfrak{D}_{K}=\prod_{\mathfrak{p}}\mathfrak{p}^{w_{\mathfrak{p}}}$, where the product is taken over all prime ideals $\mathfrak{p}$ of $\mathcal{O}_{K}$ with $e_{\mathfrak{p}}>1$, and where $w_{\mathfrak{p}}\geq e_{\mathfrak{p}}-1$. Further, $|D_{K}|=N_{K}\mathfrak{D}_{K}$. This implies that for any ideal $\mathfrak{a}$ of $\mathcal{O}_{K}$,
(4.7) $\mathcal{R}_{K}^{\prime}(\mathfrak{a})\,|\,\mathcal{R}_{K}(\mathfrak{a})\,|\,D_{K}\cdot\mathcal{R}_{K}^{\prime}(\mathfrak{a}),\ \ \text{where }\mathcal{R}_{K}^{\prime}(\mathfrak{a}):=\prod_{\mathfrak{p}|\mathfrak{a}}N_{K}\mathfrak{p}.$
Conjecture 4.12 (Masser’s uniform abc conjecture over number fields, 2002).
There is a constant $C(\epsilon)>0$ depending on $\epsilon$, such that for every $\epsilon>0$ the following holds. For every number field $K$ of discriminant $D_{K}$ and every non-zero $\alpha,\beta,\gamma\in K$ with $\alpha+\beta=\gamma$, we have
$H_{K}(\alpha,\beta,\gamma)\leq C(\epsilon)^{[K:\mathbb{Q}]}\big(|D_{K}|\cdot\mathcal{R}_{K}(\mathfrak{a}^{-3}\alpha\beta\gamma)\big)^{1+\epsilon},$
where $\mathfrak{a}$ is the fractional ideal generated by $\alpha,\beta,\gamma$.
This implies the following bound for the solutions of the $S$-unit equation (4.4) $ax+by=1$ in $x,y\in\mathcal{O}_{S}^{\ast}$, where again $S$ is a finite set of places of $K$, containing the infinite places and $a,b\in K^{*}$: let $\mathcal{R}_{S}:=1$ if $S=S_{\infty}$ and $\mathcal{R}_{S}:=\prod_{i=1}^{t}N_{K}\mathfrak{p}_{i}^{e_{\mathfrak{p}_{i}}}$, and put $\mathcal{R}_{K}(a,b):=\prod_{\mathfrak{p}}N\mathfrak{p}^{e_{\mathfrak{p}}}$, where the product is taken over all $\mathfrak{p}\in M_{K}\setminus S$ such that $|a|_{\mathfrak{p}}$ and $|b|_{\mathfrak{p}}$ are not both equal to $1$. Then for every solution $x,y\in\mathcal{O}_{S}^{\ast}$ of $ax+by=1$ we have
$H_{K}(1,x,y)\leq C(\epsilon)^{d}\big(|D_{K}|\cdot\mathcal{R}_{S}\cdot\mathcal{R}_{K}(a,b)\big)^{1+\epsilon}H_{K}(1,a,b)^{2}.$
See also Győry (2022), Theorem 3.
Some alternative effective methods were also developed to obtain effective bounds for the solutions of $S$-unit equations. Bombieri (1993, 2002) and Bombieri and Cohen (1997, 2003) worked out such an effective method in Diophantine approximation, based on an extended version of the Thue–Siegel principle, the Dyson Lemma and some geometry of numbers. Bugeaud (1998), following their approach and combining it with estimates for linear forms in logarithms, proved results which are in certain parameters sharper than those of Bombieri and Cohen.
During 1983–95 Frey initiated and developed in several papers the modular degree approach for $S$-unit equations over $\mathbb{Q}$; see e.g. Frey (1997) where he gives height bounds which became unconditional around 2000 when modularity was proved. As is surveyed by von Känel (2024), effective bounds over $\mathbb{Q}$ were proved in 2011 independently and simultaneously by von Känel (2013, 2014b) and by Murty and Pasten (2013), Pasten (2014).
However, it should be remarked that for most applications of $S$-unit equations, including the reduction theory of integral polynomials treated in our paper, more general results concerning $S$-unit equations of the form (4.4) over arbitrary number fields are needed.
4.5. A brief sketch of the proof of a less precise version of Theorem 4.2
a
Consider a monic polynomial $f\in\mathbb{Z}[X]$ of degree $n$ and discriminant $D\neq 0$. In view of Theorem 2.3 we may assume that $n\geq 3$.
First we sketch the proof of assertion (i). Assume that $f$ is irreducible over $\mathbb{Q}$. Let $K=\mathbb{Q}(\alpha)$ for a zero $\alpha$ of $f$, and denote by $D_{K}$ the discriminant of $K$. Then combining the Minkowski inequality with the fact that $D_{K}$ divides $D(f)$, i.e. $D$, (i) follows with an appropriate $c_{1}$. If now $f$ is reducible and $f=f_{1}\cdots f_{t}$ with monic irreducible $f_{1},\ldots,f_{t}$, then using $D(f_{j})\mid D(f)$ in $\mathbb{Z}$ and applying the just proved (i) for $j=1,\ldots,t$, we obtain (i) in the general case as well.
We now sketch the proof of (ii) in Theorem 4.2. Its main steps are as follows.